OpenStudy (spareb665):

For an income over $10000, a state charges a fixed tax of $300 plus 6.5% of the income in excess of $10000. A) Calculate the income tax for a person with an income of $80000. B) If Mike paid $7508.50. What was Mike's income?

OpenStudy (mathstudent55):

Call income x. Then, income in excess of 10,000 is x - 10,000. Fixed tax is 300. Tax on excess is 6.5% of excess, or 6.5% of (x - 10,000) = 0.065(x - 10,000) Total tax, T = 300 + 0.065(x - 10,000) A) use the above formula for tax and plug in x = 80,000 B) let T = 7508.5, plug that into the eqaution and solve for x.

OpenStudy (spareb665):

so 300 + 0.065(80000-10000) =4850?

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Part B) The tax is 7508.50, that equals T You are looking for x 7508.50 = 0.065(x-10000)